OpenStudy (taylor.joe):

For what value of θ, in degrees, is sin θ = cos 58°?

OpenStudy (sshayer):

\[\sin \theta=\cos 58=\cos \left( 90-32 \right)=\sin 32\] \[\theta=32\]

OpenStudy (taylor.joe):

Thank you!

OpenStudy (sshayer):

yw